ydx + (1-3y)xdy = 3y²e³ydy

1
Expert's answer
2022-03-08T05:44:01-0500

ydx + (1-3y)xdy = 3y²"e^{3y}" dy

=> "\frac{dx}{dy} + (\frac{1}{y}-3)x = 3ye^{3y}"

This is a linear differential equation of first order

Now integrating factor is given by,

I.F. = "e^{\int (\frac{1}{y}-3)dy}=e^{\ln y - 3y} = ye^{-3y}"

Multiplying both sides by integrating factor,

"ye^{-3y}" "[\frac{dx}{dy} + (\frac{1}{y}-3)x]=" "ye^{-3y}3ye^{3y}"

=> d{x"ye^{-3y}" } = 3y²dy

=> "\int{d(xye^{-3y})}" = "\int{3y²}dy"

=> xye"^{-3y}" = "\frac{3y³}{3} + C"

=> xy "=e^{3y}(y\u00b3 + C)" , "C" is arbitrary constant.

This is the general solution of the given differential equation.